The Measure of the larger of two complementary angles is 12° more than five times the measure of the smaller angle. what are the angle measures?
Answer by person24(13)

First put it in a equation. Since complementary means two angles that add up to 90 degrees. a+b=90 degrees.
We know that larger angle (b) is 5 times lager than A.
So write
5A.
We also know that is added by 12.

5A+12=B.
Now lets subsitue b for 5A+12.....
5A+12+A=90 degress
Solve for A
6A+12=90
Subtract 12
6A=90
Divide by 6
A=15
Solve for B by plugging A into the original equation A+B=90
15+B=90
Subtract 15
B=75
Hope that helps!